# Retain or Reframe? A Computational Framework for the Analysis of Framing in News Articles and Reader Comments

### v1

@@ 7.3 How Do Readers Reframe?
Table 5: Five most common reframings of the topic ‘Immigration’ in SOCC (top) and NYT (bottom) as ‘dominant article frame $\rightarrow$ dominant comment frame’.
@@ 8 Discussion
Unlike previous computational approaches to framing detection and analysis, our study is centered on a constructionist understanding of framing as an interactive process where readers selectively engage, modify, and reconstruct media frames based on their own interpretative resources  (Neuman et al., 1992; Scheufele, 1999; Gorp, 2007). While existing computational methods treat framing as isolated classification tasks focused solely on source content, we examined the dynamic interaction between media framing and audience reception. In uncovering systematicities (RQ1; Figure 4) and topic-specific differences (RQs 2-3) in a large-scale study, we bridged disconnect between large-scale yet isolated NLP approaches and small-scale communication studies that lack generalizability across topics and contexts. In applying one framing conceptualization across eleven topics, our findings suggest that inconsistent results in prior work stem from topic- and outlet-specific variation in framing effects. We also found consistent trends, such as value-laden frames (“Fairness and Equality”, “Morality”) being more likely to be reframed. This is in line with theoretical formulations of framing, according to which individuals more readily accept frames that match their pre-existing values (Chong and Druckman, 2007).
@@
Our analysis reveals outlet-specific trends — e.g., NYT readers’ political reframing and SOCC readers’ economic focus — likely reflect interactions between editorial stance, audience demographics, and national context. While frame-based patterns are relatively stable across outlets, topic effects vary more substantially. Since our comparison conflates country (US vs. Canada), editorial leaning (center-left vs. center-right), and time period (2012–2016 vs. 2017–2018), we refrain from strong interpretation. Future work should disentangle these factors to clarify framing dynamics and inform broader interdisciplinary research.
@@
A key methodological contribution of our framework lies in demonstrating reliable frame classification across fundamentally different text genres – from formal news articles to informal reader comments. This cross-genre generalization addresses a critical gap in computational framing analysis, where models typically operate within homogeneous text types. Our validation shows that sentence-level frame classification can successfully detect frames in structurally and stylistically different content, with strong performance on our human-annotated data set FrAC.
@@
Our study builds on the MFC, a widely-used framework in NLP. Yet, it is by no means the only or the optimal framework, and it has been criticized for the broad nature and lack of theoretical grounding of its frames (Ali and Hassan, 2022). We note that our method is framework-agnostic: alternative framing theories could be readily integrated, whether more generic frameworks (e.g., Semetko and Valkenburg (2000)), or issue-specific taxonomies tailored to particular domains, a rich direction for future work. Similarly, our methods for dominant frame detection and sentence-level annotation, while supporting discovery of major trends, limit finer-grained framing analyses, which could be explored in future work.
@@
Finally, our framework holds potential relevance for other NLP tasks including argumentation interpretation and generation, and even disinformation detection. The systematic patterns we identify in how audiences reconstruct frames could inform models of persuasive text generation, helping systems understand how different framings might be received and transformed by target audiences.
@@ 9 Conclusion
We introduced a scalable computational framework for analyzing media framing effects as an interactive process between source content and audience responses, grounded in a constructionist view on framing. Through analysis of over 3,000 news articles and 380,000 comments across two outlets and eleven topics, we demonstrate that readers reframe content at least as frequently as they retain original article frames. Frame retention varies systematically by frame type, with value-laden frames being more likely to be reframed, and when reframing occurs, readers consistently shift toward political interpretation regardless of the original framing. To support our analysis, we contributed a robust framing scheme from an established taxonomy, together with robust classifiers that work across different document types, providing a foundation for future research on media framing effects.
@@ Appendix E Annotation Instructions and Interface
<Displays list of framing categories as in Table 8>
@@
Table 8: Framing Categories used in this work.
@@
Appendix F Reframing Patterns in SOCC
Figure 6: Reframing patterns in the SFU Opinion Corpus (SOCC) showing the flow from primary article frames (left) to primary comment frames (right). We print the original % of frames in articles and comments, respectively. This is reflected on the bar height on the right, but not on the left to reduce visual clutter.
Appendix G Reframing Patterns by Topic
Topic SOCC NYT
Reframing Pattern Count Reframing Pattern Count
Climate Change Political and Policies → Economic 1126 Political and Policies → Economic 1590
Economic → Political and Policies 907 Economic → Political and Policies 857
Political and Policies → Cultural Identity 185 Political and Policies → Morality 690
Political and Policies → Public Opinion 181 Political and Policies → Health and Safety 686
Economic → Cultural Identity 156 Political and Policies → Cultural Identity 465
Immigration Cultural Identity → Political and Policies 134 Legality and Crime → Political and Policies 1388
Political and Policies → Cultural Identity 131 Political and Policies → Economic 1311
Health and Safety → Cultural Identity 120 Political and Policies → Legality and Crime 1139
Economic → Political and Policies 108 Legality and Crime → Economic 1024
Health and Safety → Political and Policies 99 Political and Policies → Cultural Identity 950
Syria & IS Security and Defense → Political and Policies 405 Security and Defense → Political and Policies 321
Political and Policies → Security and Defense 360 Security and Defense → Morality 165
Security and Defense → Morality 127 Health and Safety → Political and Policies 149
Security and Defense → Cultural Identity 120 Health and Safety → Morality 123
Political and Policies → Morality 96 Security and Defense → Cultural Identity 106
Trump & Elections Political and Policies → Cultural Identity 407 Political and Policies → Morality 2516
Political and Policies → Public Opinion 293 Political and Policies → Cultural Identity 2373
Political and Policies → Economic 206 Political and Policies → Health and Safety 1362
Political and Policies → Fairness and Equality 178 Cultural Identity → Political and Policies 1145
Cultural Identity → Political and Policies 145 Political and Policies → Economic 1039
Table 9: Top 5 reframing patterns by topic for SOCC and NYT outlets, showing the transition from one frame to another and the frequency count.

### v2

@@ 7.3 If readers reject the dominant article frame, how do they reframe?
Table 7: Reframing “Immigration” in SOCC (top) and NYT (bottom). Each cell shows the five most common selective (left) and complete (right) shifts from dominant article $\rightarrow$ comment frame. Frame names are abbreviated to their first word.
@@ 8 Discussion
Unlike previous computational approaches to framing detection and analysis, our study is centered on a constructionist understanding of framing as an interactive process where readers selectively engage, modify, and reconstruct media frames based on their own interpretative resources  (Neuman et al., 1992; Scheufele, 1999; Gorp, 2007). While existing computational methods treat framing as isolated classification tasks focused solely on source content, we examined the dynamic interaction between media framing and audience reception. In uncovering systematicities (RQ1; Figure˜2) and topic-specific differences (RQs 2-3) in a large-scale study, we bridged the disconnect between large-scale yet isolated NLP approaches and small-scale communication studies that lack generalizability.
In applying one unified framing conceptualization across eleven topics, our findings suggest that inconsistent results on reframing patterns in prior work can be partially explained by topic- and outlet-specific variation in framing effects. We also identified consistent trends, such as value-laden frames (“Fairness and Equality”, “Morality”) being more likely to be reframed (Figure 2), but also more likely to be introduced in the context of complete reframing (Figure 4(a)). This divergent pattern suggests that audiences reject value-laden framings that conflict with their worldview while actively introducing those that align with it. This extends theoretical formulations about audience personal formulation of messages (Chong and Druckman, 2007) by showing that readers not only selectively accept frames matching their values, but also actively impose their preferred value frames onto discussions where they were originally absent.
In distinguishing selective from complete reframing, we show that readers predominantly reframe in selective ways, providing empirical evidence for what framing theory has long posited: that news consumers reformulate and engage with content selectively based on their preferences (Neuman et al., 1992; Scheufele, 1999) – a phenomenon that small-scale social science studies could not disentangle systematically, and that NLP studies have overlooked by disregarding response analysis.
@@
Our analysis reveals that frame-based patterns are relatively stable across outlets, while topic-specific effects vary more widely. These differences likely arise from complex interactions between unobservable factors including reader demographics and national context. Furthermore, our data sets conflate country (US versus Canada), editorial leaning (center-left NYT versus center-right SOCC), and time period (2012–2016 NYT versus 2017–2018 SOCC). We hence deliberately maintain a descriptive level and leave disentanglement of these factors for future work.
@@
A key methodological contribution lies in demonstrating reliable frame classification across fundamentally different text genres – formal news articles and informal reader comments. This cross-genre generalization addresses a critical gap in computational framing analysis, where models typically operate within homogeneous text types. Our validation shows strong performance on our human-annotated data set FrAC.
@@
Our study builds on the MFC, a widely-used framework in NLP. Yet, it is by no means the only and optimal frameworks, and it has been criticized for the broad nature and lack of theoricatical ground of its framings (Ali and Hassan, 2022). We note that our method is framework-agnostic: alternative framing theories could be readily integrated, whether more generic frameworks (e.g., Semetko and Valkenburg (2000)), or issue-specific taxonomies tailored to particular domains. Similarly, our approach supports both large-scale analyses of framing trends and more fine-grained investigations that future work could further explore the role of minority frames.
@@
Finally, our framework holds potential relevance for other NLP tasks including argumentation interpretation and generation, and disinformation detection. The systematic patterns we identify in how audiences reconstruct frames could inform models of persuasive text generation, helping systems understand how different framings might be received and transformed by target audiences.
@@ 9 Conclusion
This work addresses the gap between framing theory and its computational operationalization. While framing theory posits that individuals respond differentially to framed information – accepting, rejecting, or modifying frames based on their cognitive schemas – computational approaches have largely reduced framing to taxonomic classification, neglecting audience reception processes.
We introduced a scalable computational framework for analyzing media framing effects as an interactive process between source content and audience responses. Through analysis of over 3,000 news articles and 380,000 comments across two outlets and eleven topics, we demonstrate that readers reframe content at least as frequently as they retain original frames, often reframing selectively by picking up on secondary article frames. Frame retention varies systematically by frame type with objective frames (e.g., “Health & Safety”) more likely to be retained than value-laden frames (e.g., “Morality”). We contributed a robust framing scheme with classifiers that work across document types, providing a foundation for future research.
Our analysis leaves open questions for future work, including how framing effects propagate through nested comment conversations and identifying causal linguistic properties that explain specific framing effects, or thread and temporal analysis.
@@ Appendix E Annotation Instructions and Interface
See Table 10
@@
Table 10: Framing categories shown on the annotation interface.
@@
Appendix F Reframing Patterns by Topic
Figure 11 and Figure 12 show the 5 most frequent selective and complete reframing patterns in both news outlets in coverage of climate change, immigration, Syria & IS, Trump & elections.
Topic Outlet Selective Reframing Pattern Count
$C$ SOCC Political $\to$ Economic 1001
Economic $\to$ Political 854
Political $\to$ Public Opinion 113
Economic $\to$ Cultural 90
Political $\to$ Cultural 79
NYT Political $\to$ Economic 1585
Economic $\to$ Political 857
Political $\to$ Health 428
Political $\to$ Cultural 345
Political $\to$ Legality 232
IM SOCC Cultural $\to$ Political 132
Health $\to$ Cultural 113
Political $\to$ Cultural 99
Health $\to$ Political 96
Economic $\to$ Political 85
NYT Legality $\to$ Political 1388
Legality $\to$ Economic 955
Political $\to$ Economic 927
Political $\to$ Legality 903
Political $\to$ Cultural 769
SY SOCC Political $\to$ Security 352
Security $\to$ Political 326
Security $\to$ Morality 78
Security $\to$ Cultural 64
Political $\to$ Morality 41
NYT Security $\to$ Political 321
Health $\to$ Political 146
Health $\to$ Morality 123
Security $\to$ Morality 117
Security $\to$ Cultural 79
TR SOCC Political $\to$ Cultural 319
Political $\to$ Public Opinion 256
Political $\to$ Economic 146
Cultural $\to$ Political 145
Political $\to$ Fairness 108
NYT Political $\to$ Cultural 2136
Political $\to$ Morality 2070
Cultural $\to$ Political 1145
Political $\to$ Health 983
Political $\to$ Economic 774
Table 11: Top 5 selective reframing patterns by topic and outlet. $C$ = Climate Change, IM = Immigration, SY = Syria, TR = Trump.
Topic Outlet Complete Reframing Pattern Count
$C$ SOCC Political $\to$ Economic 125
Political $\to$ Cultural 106
Political $\to$ Health 72
Political $\to$ Public Opinion 68
Economic $\to$ Cultural 66
NYT Political $\to$ Morality 473
Political $\to$ Health 258
Political $\to$ Cultural 120
Health $\to$ Morality 62
Economic $\to$ Morality 51
IM SOCC Political $\to$ Cultural 32
Economic $\to$ Political 23
Health $\to$ Morality 22
Economic $\to$ Cultural 17
Political $\to$ Fairness 17
NYT Political $\to$ Economic 384
Political $\to$ Legality 236
Legality $\to$ Cultural 189
Political $\to$ Cultural 181
Political $\to$ Morality 177
SY SOCC Security $\to$ Political 79
Security $\to$ Cultural 56
Political $\to$ Morality 55
Security $\to$ Morality 49
Political $\to$ Cultural 39
NYT Security $\to$ Morality 48
Security $\to$ Cultural 27
Security $\to$ Economic 25
Fairness $\to$ Morality 14
Health $\to$ Cultural 14
TR SOCC Political $\to$ Cultural 88
Political $\to$ Fairness 70
Political $\to$ Economic 60
Political $\to$ Legality 60
Political $\to$ Morality 38
NYT Political $\to$ Morality 446
Political $\to$ Health 379
Political $\to$ Economic 265
Political $\to$ Legality 252
Political $\to$ Cultural 237
Table 12: Top 5 complete reframing patterns by topic and outlet. $C$ = Climate Change, IM = Immigration, SY = Syria, TR = Trump.
Appendix G Reframing in SOCC
Figure 6 and Figure 7 show complete and selective reframing patterns in SOCC, for comparison against NYT results.
Figure 6: Complete reframing in SOCC showing cases where readers use frames never appearing in articles.
Figure 7: Selective reframing in SOCC showing cases where readers emphasize secondary frames differing from dominant article frames but existing elsewhere in articles.